OpenStudy (anonymous):

Solve (x + 3)(x – 3) = 40. Identify any double roots. A. {3, –3} B. {7, –7} C. {0, 3, –3} D. {0, 7 (double root)} Is it A or C and why?

OpenStudy (radar):

Take the following steps to solve this: Step 1. Multiply and clear the parentheses. It is a special product and you could do it in your head.  Getting x^2 - 9 = 40 Simplify further getting x^2=40 + 9 = 49 Step 2 solve for x \[x = \sqrt{49}\]Taking the square root of both sides.